A uniform chain of mass m and length l overhangs a table with its two third part on the table.find the work to be done by a person to put the hanging part back on the table.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
2/3rd part is on the table and 1/3rd is hanging.
Therefore the mass corresponding the hanging length is:
m mass  for  l length
than for l/3 length mass would be:
m/3.
 
Hence potential energy would be = (m/3)g(l/3) = > mgl/9 J
Hence the Work done = mgl/9.